OpenStudy (anonymous):

squareroot of -50 simplified

OpenStudy (anonymous):

\[\sqrt{-50}\]

OpenStudy (anonymous):

sqrt (-50) can be written as sqrt(50 x -1) can you finish it? use the fact that sqrt (-1) = i

OpenStudy (anonymous):

wow..i know that part its the part after that confuses me

OpenStudy (anonymous):

Ok, so sqrt(-50) = sqrt(50 x -1) = i x sqrt(50) What square number goes into 50?

OpenStudy (anonymous):

10 and 5

OpenStudy (anonymous):

They are factors of 50, but they are not square numbers as their roots are irrational. What other factors go into 50?

OpenStudy (anonymous):

2 ,5,1,10,..its all confusing me

OpenStudy (anonymous):

Ok do it in pairs 1 and 50 are factors of 50 2 and 25 are factors of 50 5 and 10 are factors of 50 which one of these 6 numbers has a nice and easy square root?

OpenStudy (anonymous):

5 and 10 or 1 and 50

OpenStudy (anonymous):

I don't think you understand me :) I mean which of these 6: 1,2,5,10,25 has a square root you can do Think of it like this, is 2^2 = 4 << nope 4 is not in our list 3^2 = 9 << nope 9 is not either 4^2 = 16 nope 5^2 = ???

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Yeah! So we can write i x sqrt (50) as i x sqrt (25 x 2) Now can you finish it? :)

OpenStudy (anonymous):

yes..thanks so i think the answer is 5i squareroot 2 ?

OpenStudy (anonymous):

Yeah, that's right! Just practice a few of them you'll get the hang of it :)

OpenStudy (anonymous):

thanks

OpenStudy (08surya):